# To Study Properties of a Known Procedure in Adaptive Sequential Sampling Design

### v2

@@ Abstract
We consider the procedure proposed by Bhandari et al., (2009) in the context of two-treatment clinical trials, with the objective of minimizing the applications of the less effective drug to the least number of patients. Our focus is on an adaptive sequential procedure that is both simple and intuitive. Our findings show that the number of applications of the less effective drug is a finite random variable whose all moments are also finite. In contrast, Bhandari et al., (2009) observed that this number increases logarithmically with the total sample size. We attribute this discrepancy to differences in their choice of starting sample size and the method of analysis employed.
@@ 1 Introduction
The field of adaptive sequential design has a long history of research, particularly in optimizing multiple objective functions. Berry and Fristedt, (1985) in their book on Bandit Problems, explored Bayesian methods to achieve optimal designs in this context.
@@
In this context, Bhandari et al., (2007) considered the case with known parameters and concluded that the expected number of applications of the less effective drug is finite under an adaptive sequential design. However, where the parameters are unknown, Bhandari et al., (2009) found that the expected number of applications of the less effective drug increases logarithmically with the total sample size. This result was derived under the assumption of a large initial sample size.
In the present work, we modify the procedure proposed in Bhandari et al., (2009) and provide a more refined analysis to elucidate that the expected number of applications of the less effective drug remains finite. In fact, we show that all the moments of the random variable that denote the number of applications of the less effective drug, are finite.
The paper is structured as follows. We explicate the procedure in Section 2, followed by a theorem with some remarks in Section 3. The proof of the theorem is provided in Section 4. In Section 5, we assess the performance of the procedure under various parameter choices using simulations. Finally, Section 6 concludes.
@@ 2 Preliminaries
The procedure described above was extensively studied by Bhandari et al., (2009), where it was referred to as Procedure III.
@@ 3 Main Result
We have the following:
@@
Bhandari et al., (2009) found that both $\frac{N^{\prime}_{1,N}}{\log(N)}$ and $\frac{\mathbb{E}(N^{\prime}_{1,N})}{\log(N)}$ tend to positive constants, contrary to our main result. This variation arises because they started with a large initial sample size and employed a different method of (mathematical) analysis.
@@ 4 Proof of Main Result
The following two results, as stated in Chapter 1, Section 9 of Billingsley, (1995), are presented here without proof.
@@
Note that, along with normal distribution for $\{X_{n}\}$, Chernoff’s theorem also holds for any distribution with $\mathbb{E}(X_{n})<0$ and $\mathbb{P}(X_{n}>0)>0$ (see Billingsley, (1995)).
Using Result 4.1, we find that for any given $\varepsilon>0$, there exists a natural number $N_{\varepsilon}^{*}$ and subset of sample space $A$ with $\mathbb{P}(A)>1-\varepsilon$, such that
@@
$M_{u}^{*}=\inf\left\{k:\sup\limits_{n\geq k}\bar{Z}_{n}\leq u\right\},$
@@
5 Simulation
In this section, we present simulation results to examine the limiting behavior of $N_{1,N}$ as the total sample size $N$ increases and to support our main result in Section 3. We implement the allocation rule described in Section 2 for various values of $(\theta_{0},\theta_{1},\sigma_{0},\sigma_{1})$ in the context of Normal populations. For each $N$, we conduct $10,000$ replications and use the simulated data to compute the probability of correct selection (PCS) as the proportion of correct decisions in these trials. In addition, we estimate $\mathbb{E}(N_{1,N})$ for different parameter settings and analyze its limiting values as $N$ increases.
Table 1: Simulation results for distinct Normal mean pairs $(\theta_{0},\theta_{1})$ and variance pairs $(\sigma_{0}^{2},\sigma_{1}^{2})=(1,0.7)$ under a fixed initial sample size.
@@ 5 Simulation
The simulation results for different mean pairs $(\theta_{0},\theta_{1})$ and initial sample sizes are summarized in Table 1. We simulate PCS values and $\mathbb{E}(N_{1,N})$ for several choices of $N$, ranging from moderate to large, across three sets of parameter values $(\theta_{0},\sigma_{0}^{2})$ and $(\theta_{1},\sigma_{1}^{2})$. The analysis indicates that $\mathbb{E}(N_{1,N})$ stabilizes to a constant as $N$ increases. Furthermore, we observe that the allocation rule assigns more samples to the population with the higher mean value. Each simulation starts with equal initial sample sizes for both populations, carefully chosen on the basis of the parameter settings.
Table 2: Simulation results for distinct Bernoulli probability pairs $(p_{0},p_{1})$ under a fixed initial sample size.
@@
Chernoff’s theorem also holds for other distributions, including Bernoulli variables. In a similar manner, we conduct simulations for Bernoulli populations, performing $10,000$ replications for three specific sets of values of parameters $(p_{0},p_{1})$, where $p_{i}$ denotes the probability of success for $i=0,1$. The corresponding simulation results are presented in Table 2.
The findings indicate that $\mathbb{E}(N_{1,N})$ tends to a constant as $N$ grows large. For each set of parameter values, we begin with equal initial sample sizes for both populations, ensuring a balanced allocation at the start of the procedure.

### v3

@@ Abstract
We consider the procedure proposed by Bhandari et al. (2009) in the context of two-treatment clinical trials, with the objective of minimizing the applications of the less effective drug to the least number of patients. Our focus is on an adaptive sequential procedure that is both simple and intuitive. Through a refined theoretical analysis, we establish that the number of applications of the less effective drug is a finite random variable whose all moments are also finite. In contrast, Bhandari et al. (2009) observed that this number increases logarithmically with the total sample size. We attribute this discrepancy to differences in their choice of the initial sample size and the method of analysis employed. We further extend the allocation rule to multi-treatment setup and derive analogous finiteness results, reinforcing the generalizability of our findings. Extensive simulation studies and real-data analyses support theoretical developments, showing stabilization in allocation and reduced patient exposure to inferior treatments as the total sample size grows. These results enhance the long-term ethical strength of the proposed adaptive allocation strategy.
@@ 1 Introduction
The field of adaptive sequential design has a long history of research, particularly in optimizing multiple objective functions. Berry and Fristedt (1985) in their book on Bandit Problems, explored Bayesian methods to achieve optimal designs in this context. There has been substantial research in the field of adaptive sequential analysis (see e.g., Friedman et al. (2010); Ivanova et al. (2000); Rosenberger et al. (2004)).
@@
Modern methodologies align well with the objectives of adaptive sequential sampling by developing advanced designs across various settings, such as covariate adjustment, ordinal responses, crossover trials and multi-treatment response adaptive design — that primarily emphasize allocation strategies (or rules) and efficiency rather than bounding the usage of suboptimal treatments (Das et al. (2023); Bandyopadhyay et al. (2020); Biswas et al. (2020); Biswas et al. (2008), pp. 3333333353535353). Das () and Das () devoted much effort to designing ethically strong procedures with reduction in allocations to the less effective treatment, including considerations for misclassifications and adaptive interim decisions. However, our approach directly focuses on this ethical aspect by ensuring finiteness of the expected number of applications of the less effective treatment.
In this context, Bhandari et al. (2007) considered the case with known face value of the parameters and concluded that the expected number of applications of the less effective drug is finite under an adaptive sequential design. However, where the parameters are totally unknown, Bhandari et al. (2009) found that the expected number of applications of the less effective drug increases logarithmically with the total sample size. This result was derived under the assumption of a large initial sample size.
In the present work, we modify the procedure proposed in Bhandari et al. (2009) and provide a more refined analysis to elucidate that the expected number of applications of the less effective drug remains finite. In fact, we show that all the moments of the random variable that denotes the number of applications of the less effective drug, are finite.
The paper is structured as follows. We explicate the two-treatment procedure in Section 2, followed by a theorem with some remarks in Section 3. The proof of the theorem is provided in Section 4. In Section 5, we introduce the extended version of the two-treatment procedure in the context of multi-treatment procedure and present its theoretical properties. Section 6 assesses the results of comprehensive simulation studies under various parameter choices. Section 7 illustrates the effectiveness of the proposed methods through real-data analyses based on two clinical trial datasets. Finally, Section 8 concludes the paper.
@@ 2 Preliminaries
The procedure described above was extensively studied by Bhandari et al. (2009), where it was referred to as Procedure III.
@@ 3 Main Result
Under the notations and assumptions given in the Section 2, we have the following theorem:
@@
Bhandari et al. (2009) found that both $\frac{N^{\prime}_{1,N}}{\log(N)}$ and $\frac{\mathbb{E}(N^{\prime}_{1,N})}{\log(N)}$ tend to positive constants, contrary to our main result. This variation arises because they started with a large initial sample size and employed a different method of (mathematical) analysis.
@@ 4 Proof of Main Result
The following two results, as stated in Chapter $1$, Section $9$ of Billingsley (1995), are presented here without proof.
@@
Note that, along with normal distribution for $\{X_{n}\}$, Chernoff’s theorem also holds for any distribution with $\mathbb{E}(X_{n})<0$ and $\mathbb{P}(X_{n}>0)>0$ (see, Billingsley (1995)).
Using Result 4.1, we find that for any given $\varepsilon>0$, there exists a natural number $N_{\varepsilon}^{*}$ and subset $A$ of sample space with $\mathbb{P}(A)>1-\varepsilon$, such that
@@
$M_{u}^{*}=\inf\left\{k:\sup\limits_{n\geq k}\bar{Z}_{n}\leq u\right\},$ (1)
@@
5 Extension to Multiple Treatments
5.1 Multi-Treatment Allocation Rule
Suppose that we have $m$ ($\geq 2$) competing drugs. Consider $m$ populations: $\mathcal{X}_{1}$, $\mathcal{X}_{2}$, …, $\mathcal{X}_{m}$. The random variables in population $\mathcal{X}_{j}$ ($1\leq j\leq m$) have i.i.d. $\operatorname{N}(\theta_{j},\sigma_{j}^{2})$ distribution. We draw samples adaptively, i.e., we draw samples sequentially and at stage $n$, after drawing a total of $n$ samples, we define the following count variables:
for $\text{for }1\leq j\leq m\text{, }N^{\prime}_{j,n}:=\text{ number of samples drawn from }\mathcal{X}_{j}\text{.}$
Let $\hat{\theta}_{n,j}$ denotes the sample mean at stage $n$ for $\mathcal{X}_{j}$ for $1\leq j\leq m$. We adopt the following allocation rule:
  1. (i)
    Increase $N^{\prime}_{j^{*},n}$ by $1$, where $j^{*}$ = $\underset{1\leq j\leq m}{\arg\max}$ $\hat{\theta}_{n,j}$,
  2. (ii)
    If there are $s$ many argmaxes, then increase those count variables with probability $\frac{1}{s}$ each.
Remark 8.
We assume $\theta_{1}>\theta_{2}>\ldots>\theta_{m}$ without any loss of generality and $\sigma_{j}^{2}$ to be unity for each $j$.
5.2 Theoretical Properties
Let us define the following:
[…]
Theorem 5.1.
Under the notations and assumptions given in the Section 5.1, we have the following theorem:
  1. (i)
  2. (ii)
    All the moments of $N_{1,N}$ are bounded (uniformly over $N$).
5.3 Proof of Theorem 5.1
Using Result 4.1, we find that for any given $\varepsilon>0$, there exists a natural number $N_{\varepsilon}^{*}$ and subset $A$ of sample space with $\mathbb{P}(A)>1-\varepsilon$, such that
$\forall\hskip 5.69054pt\omega\in A,\quad\forall\hskip 5.69054ptn\geq N_{\varepsilon}^{*},\quad\forall\hskip 5.69054ptj\in{1,\dots,m},\quad|\hat{\theta}_{j,(n)}(\omega)-\theta_{j}|<\varepsilon,$
where $\hat{\theta}_{j,(n)}(\omega)$ denotes the mean of first $n$ observations coming from the data streams $\mathcal{X}_{j}$.
Using the above, we find that with probability $\geq(1-\varepsilon)$,
This proves part (i) of Theorem 5.1.
We define $M_{u}^{*}$ as in Equation 1.
Remark 9.
[…]
Then
$N_{1,n}=\max\limits_{\begin{subarray}{c}1\leq j\leq m\\ j\neq\hat{j}_{n}\end{subarray}}\hskip 2.84526ptN^{\prime}_{j,n}\leq\sum\limits_{\begin{subarray}{c}1\leq j\leq m\\ j\neq\hat{j}_{n}\end{subarray}}\hskip 2.84526ptN^{\prime}_{j,n}\leq\sum\limits_{\begin{subarray}{c}1\leq j\leq m\\ j\neq\hat{j}_{n}\end{subarray}}\hskip 2.84526ptM_{uj}^{*}\leq\sum\limits_{1\leq j\leq m}\hskip 2.84526ptM_{uj}^{*}.$
Therefore, $N_{1,n}\leq\sum\limits_{1\leq j\leq m}\hskip 2.84526ptM_{uj}^{*}$.
Each $M_{uj}^{*}$ has bounded moments (using Lemma 4.1). Also, $M_{uj_{1}}^{*}$ and $M_{uj_{2}}^{*}$ are independent $\forall\hskip 2.84526ptj_{1}\neq j_{2}$, as the populations are independent. So, $N_{1,n}$ also has finite moments.
This proves part (ii) of Theorem 5.1.
6 Simulation Study
We conduct simulation studies to evaluate the performance of our adaptive sampling procedure as outlined in Section 2. In each replication, we generate samples of $X_{i}$’s and $Y_{j}$’s adaptively according to the allocation rule and the estimators $\hat{\theta}_{0,n}$ and $\hat{\theta}_{1,n}$ are updated sequentially at each stage $n$. The procedure continues until the stopping condition is met at $n=N$, upon which we record the values of $N_{1,N}$ and whether a correct selection (CS) was made, indicated by a binary outcome ($1$ for correct, 00 for incorrect).
This entire procedure is repeated $10,000$ times for each setting, and we compute the empirical average of $N_{1,N}$ and the proportion of CS to estimate $\mathbb{E}(N_{1,N})$ and the probability of correct selection (PCS), respectively.
To examine the limiting behavior of $N_{1,N}$ as the total sample size $N$ increases and to support the theoretical findings of Section 3, we implement the allocation rule described in Section 2 for various parameter combinations $(\theta_{0},\theta_{1},\sigma_{0},\sigma_{1})$ in the context of Normal populations. For each $N$ (from moderate to large), we conduct $10,000$ replications and use the simulated data to compute PCS as the proportion of correct decisions. In addition, we estimate $\mathbb{E}(N_{1,N})$ for different parameter settings and analyze its limiting values as $N$ increases.
The simulation results for different mean pairs $(\theta_{0},\theta_{1})$ are summarized in Table 1. The analysis indicates that $\mathbb{E}(N_{1,N})$ stabilizes to a constant as $N$ increases. The allocation rule consistently assigns more samples to the population with a higher mean value. Each simulation starts with equal initial sample sizes for both populations, determined on the basis of the underlying parameters.
Table 1: Simulation results for two Normal populations with distinct mean pairs $(\theta_{0},\theta_{1})$ and constant variance pair $(\sigma_{0}^{2},\sigma_{1}^{2})=(1,0.7)$ under a fixed initial sample size for two-treatment procedure.
@@ 6 Simulation Study
Chernoff’s theorem also holds for other distributions. To demonstrate the broader applicability of our approach, we conduct a similar simulation study for Bernoulli populations. The simulation results for three specific parameter configurations of $(p_{0},p_{1})$, where $p_{i}$ denotes the success probability for population $i=0,1$, are presented in Table 2. As in the case of Normal distributions, the findings indicate that $\mathbb{E}(N_{1,N})$ converges to a finite value as $N$ increase. Each scenario begins with equal initial sample sizes for both populations to ensure a balanced start of the procedure.
Table 2: Simulation results for two Bernoulli populations with distinct success probability pairs $(p_{0},p_{1})$ under a fixed initial sample size for two-treatment procedure.
@@
In addition, we analyze cases where both populations are identically distributed. Table 3 reports simulation results for this setting under both Normal and Bernoulli populations. These represent scenarios in which the treatments are equally effective. Accordingly, PCS is approximately $0.5$, reflecting the indistinguishability of the two treatments. Notably, $\mathbb{E}(N_{1,N})$ increases slowly but remains proportionally smaller relative to $N$, underscoring the bounded nature of the allocation count even under treatment equivalence.
Table 3: Simulation results based on two identically distributed populations under a fixed initial sample size for two-treatment procedure.
$\operatorname{N}(1,1)$ populations
Total Sample Size $(N)$ PCS $\mathbb{E}(N_{1,N})$ $\min\{\mathbb{E}(N^{\prime}_{0,N}),\mathbb{E}(N^{\prime}_{1,N})\}$ $\min\{\mathbb{E}(N^{\prime}_{0,N}),\mathbb{E}(N^{\prime}_{1,N})\}/N$
  200 0.4987   56.5952    99.8838 0.499419
  300 0.4994   61.9217   149.6241 0.498747
  400 0.5098   65.9665   197.1139 0.492785
  800 0.5084   76.9530   394.5546 0.493193
  900 0.4994   79.9105   449.5351 0.499483
1000 0.5015   79.7329   498.0089 0.498009
1500 0.4949   88.1929   743.1995 0.495466
2000 0.4962   91.7086   993.3254 0.496663
2500 0.4939   96.0888 1237.8096 0.495124
3000 0.5069   97.8849 1477.5369 0.492512
3500 0.5009 101.4421 1746.3167 0.498948
Bernoulli($0.5$) populations
Total Sample Size $(N)$ PCS $\mathbb{E}(N_{1,N})$ $\min\{\mathbb{E}(N^{\prime}_{0,N}),\mathbb{E}(N^{\prime}_{1,N})\}$ $\min\{\mathbb{E}(N^{\prime}_{0,N}),\mathbb{E}(N^{\prime}_{1,N})\}/N$
  200 0.5047   56.1546    99.5748 0.497874
  300 0.5014   61.7647   149.9355 0.499785
  400 0.5054   65.7200   198.4916 0.496229
  800 0.5016   77.5626   398.6040 0.498255
  900 0.5060   79.6830   445.1634 0.494626
1000 0.5060   79.5445   494.7277 0.494728
1500 0.5007   87.8227   749.0695 0.499380
2000 0.5048   90.6309   992.2599 0.496130
2500 0.5003   93.9105 1247.3643 0.498946
3000 0.4989   99.5797 1497.4663 0.499155
3500 0.5073 103.4383 1726.3047 0.493230
Table 4: Simulation results for Normal populations with distinct mean triplets $(\theta_{1},\theta_{2},\theta_{3})$ and constant variance triplet $(\sigma_{1}^{2},\sigma_{2}^{2},\sigma_{3}^{2})=(1,0.7,0.5)$ under a fixed initial sample size for multi-treatment procedure.
Total Sample $(\theta_{1},\theta_{2},\theta_{3})=(0.9,0.2,0)$ $(\theta_{1},\theta_{2},\theta_{3})=(2,1.2,0.5)$
[…]
        200 0.9475   9.3374 0.9527 6.9510
        300 0.9459   9.6494 0.9503 7.0553
        400 0.9492 10.0209 0.9553 7.0896
        800 0.9425 10.8458 0.9503 7.9133
        900 0.9442 11.0578 0.9550 7.4772
      1000 0.9455 11.2956 0.9577 7.7322
      1500 0.9504 11.7424 0.9515 8.0190
      2000 0.9444 12.3375 0.9561 8.1553
Table 5: Simulation results for two Normal populations with distinct mean pairs $(\theta_{0},\theta_{1})$ and constant variance pair $(\sigma_{0}^{2},\sigma_{1}^{2})=(1,0.7)$ based on Procedure III of Bhandari et al. (2009) under a fixed initial sample size.
     Size $(N)$ PCS $\mathbb{E}(N^{\prime}_{1,N})$ $\mathbb{E}(N^{\prime}_{1,N})/\log(N)$ PCS $\mathbb{E}(N^{\prime}_{1,N})$ $\mathbb{E}(N^{\prime}_{1,N})/\log(N)$ PCS $\mathbb{E}(N^{\prime}_{1,N})$ $\mathbb{E}(N^{\prime}_{1,N})/\log(N)$
        200 0.9436   24.6724   4.6566 0.9725   19.8617 3.7487 0.9442   24.7791   4.6768
        300 0.9408   31.3563   5.4975 0.9725   22.6578 3.9724 0.9455   29.9178   5.2453
        400 0.9404   37.3903   6.2406 0.9699   26.3717 4.4015 0.9447   35.9691   6.0034
        800 0.9401   61.8690   9.2554 0.9756   34.0681 5.0965 0.9415   60.8365   9.1010
        900 0.9393   68.3112 10.0422 0.9728   39.2264 5.7666 0.9432   65.5061   9.6299
      1000 0.9372   76.6053 11.0898 0.9692   45.3261 6.5616 0.9410   73.0315 10.5724
      1500 0.9377 107.6991 14.7266 0.9725   55.8550 7.6375 0.9405 103.3666 14.1342
      2000 0.9411 131.7802 17.3374 0.9738   67.7209 8.9096 0.9393 135.5299 17.8308
      2500 0.9444 153.7988 19.6572 0.9697   91.2159 11.6584 0.9354 176.4448 22.5516
      3000 0.9429 185.7885 23.2051 0.9733   95.9255 11.9812 0.9442 181.9366 22.7240
      3500 0.9415 218.7154 26.8017 0.9733 108.5402 13.3007 0.9426 214.6163 26.2993
Table 6: Simulation results for two Bernoulli populations with distinct success probability pairs $(p_{0},p_{1})$ based on Procedure III of Bhandari et al. (2009) under a fixed initial sample size.
     Size $(N)$ PCS $\mathbb{E}(N^{\prime}_{1,N})$ $\mathbb{E}(N^{\prime}_{1,N})/\log(N)$ PCS $\mathbb{E}(N^{\prime}_{1,N})$ $\mathbb{E}(N^{\prime}_{1,N})/\log(N)$ PCS $\mathbb{E}(N^{\prime}_{1,N})$ $\mathbb{E}(N^{\prime}_{1,N})/\log(N)$
        200 0.9707   20.0885   3.7915 0.9644   21.0589   3.9746 0.9693   20.1448   3.8021
        300 0.9694   23.3904   4.1009 0.9667   24.0227   4.2117 0.9707   22.9941   4.0314
        400 0.9686   26.7283   4.4611 0.9655   27.8444   4.6473 0.9720   25.2552   4.2152
        800 0.9722   36.5467   5.4673 0.9657   41.2123   6.1652 0.9723   36.5258   5.4642
        900 0.9695   41.6838   6.1278 0.9670   43.9721   6.4642 0.9710   40.0918   5.8938
      1000 0.9712   43.0519   6.2324 0.9644   49.4955   7.1652 0.9686   45.1317   6.5335
      1500 0.9691   60.6695   8.2959 0.9655   65.7791   8.9945 0.9699   59.2646   8.1038
      2000 0.9700   74.1982   9.7618 0.9608   92.1311 12.1211 0.9701   73.8442   9.7152
      2500 0.9679   94.6896 12.1024 0.9633 106.1061 13.5615 0.9692   90.5637 11.5750
      3000 0.9708 101.9465 12.7332 0.9655 117.9213 14.7284 0.9720   98.1398 12.2577
      3500 0.9711 115.4750 14.1504 0.9666 130.7154 16.0180 0.9704 117.2357 14.3662
We further extend our simulations to a multi-treatment scenario involving three competing treatments, as motivated by the theoretical framework in Section 5. In this case, we focus on the second-largest allocation count, as the maximal is invariably associated with the best-performing treatment under the allocation rule. The corresponding results, shown in Table 4, again indicate that this second-largest allocation remains finite as $N$ increases, consistent with our theoretical predictions.
For comparative purposes, we provide simulation results in Table 5 and Table 6 based on Procedure III of Bhandari et al. (2009). These demonstrate a key difference between our method and their earlier proposal: in their procedure, $\mathbb{E}(N^{\prime}_{1,N})/\log(N)$ stabilizes to a constant, whereas our approach maintains that $\mathbb{E}(N_{1,N})$ stays uniformly bounded across increasing $N$. This highlights the efficiency of our adaptive allocation strategy in the long-run performance.
Finally, we note that PCS is estimated as the mean of $10,000$ independent Bernoulli trials, each taking values 00 or $1$. Hence, by basic binomial variance arguments, the standard error of the PCS estimate is bounded above by $\sqrt{\frac{1}{4\times 10000}}=0.005$. This ensures high precision of the simulation estimates and justifies the reliability of our numerical conclusions.
Remark 10.
Note that, to compare our proposed two-treatment procedure with Procedure III of Bhandari et al. (2009), identical parameter settings were used for both the Normal and Bernoulli simulations. A comparison between Tables 1 and 2 (representing our method) and Tables 5 and 6 (representing Procedure III) yields that our analyses are far more effective under the respective parameter configurations.
7 Real Data Analysis
We consider the following datasets to illustrate the practical relevance and effectiveness of our proposed method.
7.1 Pregabalin Drug Trial for Postherpetic Neuralgia (PHN)
We analyze the performance of our proposed two-treatment adaptive allocation strategy, using data from a randomized, double-blind, placebo-controlled trial to evaluate the efficacy of the drug Pregabalin for treating postherpetic neuralgia (PHN), as described in Dworkin et al. (2003) and revisited in Biswas et al. (2008), pp. $47$.
In the original trial, patients were randomly assigned to either Pregabalin or placebo arms, and pain intensity was measured using an $11$-point numerical rating scale over an $8$-week period, where lower scores indicate better outcomes. The endpoint mean pain scores were $3.60$ (Pregabalin) and $5.29$ (placebo), with corresponding standard deviations $2.25$ and $2.20$, respectively.
As we do not have access to the raw dataset, we instead reconstruct a typical example of the data structure using these summary statistics. To adapt this continuous response setting to our model-based framework in Sections 24, we modeled the responses for Pregabalin and placebo groups as independent Normal distributions, using the reported endpoint mean scores and standard deviations. As lower mean pain scores indicate greater efficacy, the negation of the original mean scores reflects the assumption that higher values are favorable in our framework. Specifically, we use the following values as the true parameters:
  • Pregabalin (treatment 00): mean $\theta_{0}=-3.60$, sd $\sigma_{0}=2.25$,
  • Placebo (treatment $1$): mean $\theta_{1}=-5.29$, sd $\sigma_{1}=2.20$.
Using these parameters, we redesign sequential trials and generated $10,000$ replications following the two-treatment adaptive allocation rule for varying total sample sizes $N$ and computed the PCS as well as ($\mathbb{E}(N_{1,N})$), as considered in Section 6. The results are presented in Table 7.
Table 7: Results for Pregabalin trial dataset.
Total Sample Size $(N)$ PCS $\mathbb{E}(N_{1,N})$
  200 0.9381 7.4673
  300 0.9346 7.6999
  400 0.9346 7.7434
  800 0.9326 8.5777
  900 0.9380 8.4504
1000 0.9372 8.3561
1500 0.9362 8.5037
2000 0.9376 9.2875
The adaptive rule showed a marked preference for allocating patients to the superior treatment (Pregabalin), resulting in a high PCS (approximately $0.94$) and reduced average sample size for the inferior treatment (placebo) i.e., consistent with our theoretical results in Theorem 3.1. These calculations show a significant reduction in sample size allocated to the less effective treatment.
7.2 Fluoxetine drug trial for Depressive Disorder
To evaluate the adaptive performance in real clinical settings with binary response, we examined a well-documented adaptive clinical trial conducted by Tamura et al. (1994) on the antidepressant drug Fluoxetine, compared to a placebo, which employed an adaptive design using a randomized play-the-winner rule (RPTW) and further discussed in Biswas et al. (2008), pp. $46$.
In this trial, the binary response variable was defined by the Hamilton Depression Rating Scale (HAMD17): patients achieving at least a $50\%$ reduction in HAMD17 score after a minimum of $3$-weeks of therapy were labeled as responders. We apply the adaptive allocation methodology specifically to those patients assigned to the shortened REML stratum as considered in Biswas et al. (2008). To fit this into our framework introduced in Section 2, we treat the response outcomes as arising from Bernoulli distributions for both treatments — Fluoxetine and placebo. Specifically, we consider the observed sample proportions of responders for Fluoxetine and placebo as the true success probabilities $p_{0}$ and $p_{1}$, respectively. The reported response rates were:
  • Fluoxetine (treatment 00): $p_{0}=0.58$,
  • Placebo (treatment $1$): $p_{1}=0.36$.
Using these parameters, we construct a representative data structure which is used to redesign a series of sequential trials under our proposed adaptive allocation rule and stopping conditions described in Sections 24. We then carry out the trial for $10,000$ independent replications using $p_{0}=0.58$ and $p_{1}=0.36$ as the true success probabilities. Table 8 summarizes the results.
Table 8: Results for Fluoxetine trial dataset.
Total Sample Size $(N)$ PCS $\mathbb{E}(N_{1,N})$
  200 0.9419 21.5476
  300 0.9373 22.1276
  400 0.9412 22.4579
  800 0.9424 22.8271
  900 0.9384 22.9035
1000 0.9415 23.1794
1500 0.9411 23.0954
2000 0.9407 23.5513
Across increasing total sample sizes $N$, we observe a consistently high PCS (approximately $0.94$), affirming the design’s ability to identify the superior treatment. Notably, the number of patients allocated to the less effective treatment (placebo) — remained stably low, supporting our theoretical result from Section 3. Similar to the conclusion drawn from the data analysis in Section 7.1, the proposed method may also have contributed to a reduction in sample size allocated to the less effective treatment in this case also.
Remark 11.
These analyses exemplify the broad applicability and effectiveness of the proposed adaptive allocation rules under realistic clinical contexts where treatment effects are significantly different — as seen in the case of continuous responses for Pregabalin’s efficacy in managing PHN-related pain and binary responses in the Fluoxetine drug trial for depressive disorder. In contrast to prior findings by Bhandari et al. (2009), where the allocation count to the less effective treatment was shown to grow logarithmically with sample size under a large initial sample regime, our analyses show boundedness of $\mathbb{E}(N_{1,N})$ without requiring large pilot samples or complex tuning — thus confirming the practical and ethical efficiency of our framework. Both analyses emphasize the ethical advantage of our design: minimizing patient exposure to inferior treatments while maintaining statistical efficiency. This supports the effectiveness of our allocation strategy under realistic conditions in general, and in particular for medical studies, as shown in the datasets considered.